OpenStudy (anonymous):

If the area of an isosceles right triangle is 8, what is the perimeter of the triangle?

OpenStudy (anonymous):

ok that was wrong. since \[\frac{1}{2}bh =8\] \[bh=16\] \[b=h=4\]

OpenStudy (anonymous):

so both legs of the right triangle are 4 and the hypotenuse is \[\sqrt{4^2+4^2}=\sqrt{32}=4\sqrt{2}\]

OpenStudy (anonymous):

\[a. 8+\sqrt{2}\]

OpenStudy (anonymous):

add up to get the perimeter, no it is \[8+4\sqrt{2}\]

OpenStudy (anonymous):

b. \[b. 8+4\sqrt{2}\]

OpenStudy (anonymous):

yeah, that one

Directrix (directrix):

1/2 x^2 = 8 x^2 = 16 x = 4 Perimeter = 4 + 4 + 4√2 = 8 + 4√2

OpenStudy (anonymous):

A = (1/2) a^2 = 8 ( isocelses) -> a^2 = 16 => a = 4 => hyp = sqrt( 16 + 16) = 4 sqrt (2) Thus P = 4 + 4 + 4V2 = 8 + 4V2 = 13.66

OpenStudy (anonymous):

Note, this is isosceles *right triangle*, hence we can conclusively say that \( bh=16 \implies b=h=4\)